OpenStudy (anonymous):

What is the equation of the line, in general form, that passes through the point (-1, -1) and is parallel to the line whose equation is x + y = 3? A. x-y+2=0 B. x+y+2=0 C. x+y-2=0

OpenStudy (zzr0ck3r):

what is the slope of the line whose equation is \(x+y=3\)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorta

OpenStudy (zzr0ck3r):

that was not a yes or no question

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i don't know

OpenStudy (zzr0ck3r):

if we subtract x from both sides we have \(y=-x+3\)

OpenStudy (zzr0ck3r):

what is the slope of this line?

OpenStudy (anonymous):

i don't know. Kinda rushing right now

OpenStudy (zzr0ck3r):

the equation of a line is \(y=mx+b\) where m is the slope of the line, and b is the y intercept. so you have \(y=-x+3\implies y=(-1)x+3\) so what is the slope?

OpenStudy (anonymous):

-4?

OpenStudy (zzr0ck3r):

what is the number next to x? in \(y=(-1)x+3\)

OpenStudy (anonymous):

-1

OpenStudy (zzr0ck3r):

ok , so we want a line that is parallel to that, so we want a line that has the same slope, so we want a line that has a slope of -1 now our line will look like this \(\large y=mx+b\) but \(m = -1\) so \(\large y=(-1)x+b\) i.e. \(\large y = -x+b\) you with me so far? we are almost done...

OpenStudy (anonymous):

kinda... I've been doing algebra since 11pm last night so I would appreciate if I could just have the answer

OpenStudy (zzr0ck3r):

we are not here to give you answers, sorry. They get mad at us for that. we have \(y=-x+b\) we need only to find out what b is, we do this by using the poine they gave us \((-1,-1)\). plug those values in for x and y, \(y=-x+b\iff-1=-1(-1)+b\) so \(-1=1+b\) what is b?

OpenStudy (zzr0ck3r):

poine should say point....

OpenStudy (anonymous):

2

OpenStudy (zzr0ck3r):

\(-1=1+b\) subtract 1 from both sides, what is b? If you cant answer this, you might need to review some earlier stuff.